Three consecutive numbers have a sum of 132. Find the number.
This question seems like it should be so easy but, it isn't! Could you please help me?

Answer by checkley77(12569)

X+(X+1)+(X+2)=132
3X+3=132
3X=132-3
3X=129
X=129/3
X=43 ANS. FOR THE FIRST NUMBER.
43+1=44 ANS. FOR THE SECOND NUMBER.
43+2=45 ANS. FOR THE LARGEST NUMBER.
PROOF:
43+44+45=132
132=132